# Readout electronics for low occupancy High-Pressure Gas TPCs

### v1

@@ Abstract
High-Pressure Gas Time Projection Chambers have benefits such as low energy threshold, magnetisability, and acceptance, making them ideal for neutrino experiments such as DUNE. We present the design of an FPGA-based solution optimised for Gaseous Argon Near Detector (ND-GAr), which is part of the Phase-II more capable near detector for DUNE. These electronics reduce the cost significantly compared to using collider readout electronics which are typically designed for much higher occupancy and therefore, for example, need much larger numbers of FPGAs and power per channel. We demonstrate the performance of our electronics with the Teststand for an Overpressurised Argon Detector (TOAD) at Fermilab in the US at a range of pressures and gas mixtures up to , reading out channels from a multi-wire proportional chamber. The operation took place between April and July of 2024. We measure the noise characteristics of the system to be sufficiently low and we identify sources of noise that can be further mitigated in the next iteration. We also note that the cooling scheme used in the test requires improvement before full-scale deployment. Despite these necessary improvements, we show that the system can fulfil the needs of a HPgTPC for a fraction of the price of collider readout electronics.
@@ 1 Introduction
The detector’s position resolution is a critical factor in achieving the required momentum resolution and detection thresholds for complex low-energy interactions with high multiplicity. The resolution obtainable is closely related to the number of readout channels that the detector can accommodate and the sampling frequency of those channels. Whilst the numbers are not final, instrumenting a diameter TPC sufficiently to achieve the desired low energy thresholds for the DUNE would require channels with sampling frequency [1]. The development of readout electronics in the particle physics community has previously focused on the design and operation of high-throughput radiation-hard electronics required at the Large Hadron Collider (LHC)-based experiments [4]. Naively scaling such designs for neutrino experiments would be prohibitively expensive. For example, in the context of the DUNE Phase-II upgrade, adopting electronics like those used for sPHENIX [5] or the TPC of A Large Ion Collider Experiment (ALICE) [6] with its readout electronics requires a high-specification Field Programmable Gate Array (FPGA) to readout groups of 10 front-end cards [7], providing a resolution of 1- in the TPC drift direction. However, given the hundreds of thousands of channels required to meet DUNE’s resolution needs, this approach would result in a total cost larger than possible, given budgetary constraints for the DUNE near detector. In fact, the electronics were previously listed as the most expensive component for the Gaseous Argon Near Detector (ND-GAr) for DUNE.
@@
In the following sections we present the design of a scalable highly re-configurable FPGA based solution optimised to be affordable for the ND-GAr DUNE Phase-II detector [3]. This system delivers significant cost-saving compared to the systems described above while showing no performance degradation in an HPgTPC environment. A prototype has been produced and was then sent for testing to Fermilab as part of the Teststand for an Overpressurised Argon Detector (TOAD). DUNE has not decided on an amplification stage for its HPgTPC but a Gas Electron Multiplier (GEM)-based solution is the current front runner. Conversely, TOAD uses a Multi-Wire Proportional Chamber (MWPC) from the ALICE, due to its availability at the time of the test. This chamber, called an Outer Read-Out Chamber (OROC) [8], is deployed in the high-pressure vessel from the UK high-pressure facility. Large parts of this setup have previously been used to test gas mixtures for high-pressure TPCs [9].
@@ 2 Readout setup
A key challenge for future neutrino experiments is the harsh conditions in which the electronics must operate. In contrast to collider experiments, those conditions are not due to extensive radiation but instead factors such as cryogenic temperatures present in Liquid Argon (LAr)-based detectors and high pressure such as in an HPgTPC based detector. In the case of the high-pressure detectors an important consideration is the number of feedthroughs required to power and communicate with the electronics, so there is a need to aggregate data inside the vessel. Whilst power dissipation and thermal management present additional challenges, they can be addressed using water cooling through a small number of vessel penetrations. Hence, a readout system will need to optimise the number of cables required to stream data from the large amount of analogue readout channels.
@@
The proposed solution outlined in this paper uses a fast-sampling Front-End Card (FEC) with zero-suppression capability and a custom FPGA-based Power, Aggregation and Timing (PAT) card, both residing in-situ within the high-pressure volume. Data are then streamed over a standard Ethernet-based network to the DUNE Data Acquisition (DAQ) software operating on standard servers, avoiding extensive requirements within the pressurised environment. The block diagram of the proposed readout system is shown in figure 1.
@@
Compared to a collider experiment, the channels in neutrino experiments have very low occupancy (in the similarly sized ALICE, collisions occur at and each generate thousands of tracks, whereas in DUNE, beam spills happen at and generate only tens of tracks per spill). This difference means far less processing is required as there are many fewer above threshold hits. Therefore, the system can have fewer FPGAs per channel (1 per channels with the particular FECs in this design) and use much lower specification FPGAs as we stream out all hits to a software-based DAQ. Furthermore, the lower data rates from the FEC allow the native transceivers on the SAMPAs to be used to communicate with the PAT so we also do not need the additional transceiver modules that are required for higher occupancy detectors [7]. These differences result in a much lower cost to instrument the hundreds of thousands of channels necessary for a neutrino HPgTPC. Even though it is a prototype system manufactured at small scale, the readout for TOAD (including FECs, PAT cards and all the hardware to connect and power them) cost less than $ for channels. Scaling this up to the full DUNE Phase-II near detector (ND), ignoring likely economies of scale, gives a conservative cost estimate of $1.4 million.
@@ 2.1 Hardware layout
As described, our system requires two types of custom hardware boards. The first type are the FEC boards, which host the digitiser Application Specific Integrated Circuits and connect to the detector itself. The second type are the PAT boards, which utilise an FPGA device for interfacing and aggregating the data for a multitude of FECs as well as providing them with power and distributing clock signals. Both the FECs and the PATs are custom designed, with the FECs designed for the TOAD test beam detector and its ALICE OROC. The flexibility of the FPGA-based design of the PAT cards means that the PAT can be reused even if the FECs change, so, whilst designed for TOAD, this board could be used with minimal modifications for the final ND-GAr detector or other detectors with similar data volume and occupancy. In designing the hardware, we avoided components with known air gaps e.g. lidded FPGAs packages. However, as there is no standard for electronics operation in the regime, a goal of TOAD is to test the operation of components at high pressures.
@@ 2.1.1 Front-End Card
GEMs are currently the leading candidate for the charge readout for ND-GAr. The SAMPA ASIC used for this readout system was developed by the ALICE collaboration for the GEM upgrade of their TPC [7]. This upgraded detector has similar characteristics to those needed for ND-GAr, so SAMPA is expected to meet the requirements of ND-GAr. The parameters of the SAMPA measured by ALICE can be seen in table 1 [10, 7]. The electronics system described here is therefore designed around deploying the SAMPA ASIC [10] used by ALICE, in a high-pressure environment, and at a much lower cost than in the full ALICE system.
@@
Another important reason for the choice of the SAMPA chips is their ability to sample at with fast Analogue to Digital Converters. Gas detectors have about an order of magnitude faster electron drift velocities than liquid; hence, this high sample frequency is necessary.
For TOAD, the ALICE MWPC has a group of three connectors per slot, which correspond to a total of 64 channels. Given that the SAMPA ASIC has only 32 channels, the FEC designed for TOAD, which can be seen in figure 2, has three connectors and hosts two SAMPAs to capture all 64 channels. The final ND-GAr amplification stages may have a different number of channels per connector, however the more SAMPAs per FEC the lower the cost of the system so designing for 2 SAMPAs represents a reasonable worst-case cost. To minimise the required readout channels connecting to the PAT board, the two SAMPAs on each FEC are daisy-chained as the low occupancy of neutrino experiments does not require the full bandwidth of the SAMPA readout link. However, as this is a prototype detector, the layout of the FEC has been designed to allow each SAMPA to be read out individually with minimal changes to the fitted components and no re-design of the Printed Circuit Board (PCB).
For TOAD, the FECs were operated in the configuration shown in table 2. This configuration gives an expected current draw of /channel with the FEC operating at full capacity. It is not expected that the FEC will have to operate at full power due to low occupancy and zero suppression. Whilst the design of the FEC is specific to the setup for TOAD, the operation at TOAD showcases the expected performance of SAMPA in a setup at an environment much closer to that required of a HPgTPC such as ND-GAr than previously tested.
@@ 2.1.2 Power Aggregation and Timing cards
The PAT card’s primary function is to interface between the FEC cards and a standard ethernet network where data can be streamed to servers for data selection and storage. The prototype, designed at Imperial College London, utilises the advantages of FPGAs as the main data processing unit for the PAT. An image of the prototype board can be seen in figure 3. The primary components on the PAT cards are listed in table 3.
| Component | Quantity |
| Kintex Ultrascale FPGA | |
| Insulation-Displacement Contact (IDC) connectors | |
| Clock Synthesizers | |
| Direct Current to Direct Current (DC/DC) Converters |
Data aggregation is a primary requirement that drives the design of the PAT. The cards are designed to take inputs from 56 FECs each streamed over a differential pair into the FPGA package. The received data payloads from the many FECs are then extracted and propagated through a multitude of buffers and Round-Robin (RR) logic to convert them into a single output stream. This output data stream is then wrapped into an Ethernet-based protocol and sent over the FPGA serial transceivers to the COTS network switch. The full layout can be seen in figure 4. Due to the low occupancy of the detector, the aggregation factor of 56 to 1 can be achieved easily within the FPGA fabric.
As well as data aggregation, the powering and timing roles of the PAT are also crucial. For powering, we investigated two possible solutions, for which a summary of key metrics is provided in table 4. The first solution provides to each PAT via dedicated terminals on the board. The FPGA DC/DC converter then outputs which is further fed into the FEC DC/DC converter to be transformed to the needed by the FEC on-board Low-Dropout regulator (LDO). The advantage of this scheme is that each PAT card receives a dedicated power line, minimising potential interference with signal traces on the board. The disadvantage is that ND-GAr will need dedicated power feedthroughs in the pressure vessel, which increases the overall feedthrough count.
The second ‘power over serial’ powering scheme is inspired by Power over Ethernet (PoE), and aims to allow PAT cards placed within the pressure vessel to receive power over the incoming network cable. The circuit diagram used can be seen in figure 5. However, some modifications to the normal PoE standard are required due to the intended environment. PoE works by injecting a Direct Current (DC) component on top of the alternating data component using transformers. Copying the same methodology for a magnetised detector wouldn’t be viable, as transformer performance may be impacted by the surrounding magnetic field. Therefore, an alternative system was developed using air-cored inductors to extract the differential signal from the bias voltage.
@@
The implementation on an early prototype for the PAT cards was found not to give sufficient signal stability, so further optimisations are required to achieve a production level system with this powering option. This work is underway with a new dedicated hardware board allowing modifications to both the power injection and the power decoupling circuits for further analysis. However, as it has not yet converged we therefore chose the first solution (direct power) for the board design discussed here.
In addition, a study, summarised in table 4, was done to estimate the number of feedthroughs needed for both options and also a third potential solution which would use optical fibres to stream out the data and a separate power line for each PAT. This option allows several PAT cards to share one output feedthrough as optical fibres are smaller in form factor than the electrical connectors needed.
@@
Decisions on the power distribution of the system also have an impact on the timing. For the Power over Serial method, timing can be provided to the upstream PAT card that is outside the pressure vessel and all the cards within the vessel can then be synchronised by the serial transceiver recovered clock. In the case of Direct Power, the data stream is separate to power and is running through commercial Small Form-Factor Pluggable (SFP) modules. We therefore require a separate dedicated input for clock distribution. The PAT cards are designed to be able to interface with the DUNE timing system [11] over an optical fibre. In order to implement this, extra optical feedthroughs are required to receive all the timing fibres into the pressure vessel. The number of feedthroughs can be mitigated by fanning out the signal inside the vessel but there are limits to the number of boards that can be supplied from any one fibre. In both cases, the clock received at the PAT cards in the pressure vessel is transmitted from the FPGA to a dedicated clock synthesizer for cleaning and then further transmitted to the FECs.
Furthermore, distribution of the FEC clock, sync and slow control signals happens, depending on the use case, via fan-out chips or multiplexers. Given that the designed readout is synchronous to the DUNE clock, all the FEC cards are expected to receive the same copy of the clock sync signal. For this reason, an organisation of eight FECs in seven groups have been designed. The advantage of this method relies on the layout design of the PCB. The chosen FPGA can drive seven individual groups with the same clock sync signal pulse. Each of these seven individual groups is then copied in phase to all eight FECs connected to that pin. The main reason for deciding to fan-out 1:8 was to ensure that the clock quality provided to all the FECs wasn’t impacted by the fan-out ASIC. Equivalently, for the slow control interface, a series of multiplexers is implemented allowing individual access to all the FECs, providing enhanced telemetry of the FEC operation.
Finally, the PAT cards are mounted at 3 positions on the OROC, in such a way that when all 18 OROCs are mounted into a circle, the PAT cards are shared between neighbouring OROCs. In this way, the cable lengths required between FECs and PAT cards are minimised and so, these connections are optimised. The PAT cards are mounted on aluminium plates, which, through thermal contact with the Read-Out Chamber (ROC), provide some amount of passive cooling.

### v2

@@ Abstract
High-Pressure Gas Time Projection Chambers have benefits such as low energy thresholds, magnetisability, and acceptance, making them ideal for neutrino experiments such as DUNE. We present the design of an FPGA-based solution optimised for Gaseous Argon Near Detector (ND-GAr), which is part of the Phase-II more capable near detector for DUNE. These electronics reduce the cost significantly compared to using collider readout electronics, which are typically designed for much higher occupancy and therefore, for example, need much larger numbers of FPGAs and power per channel. We demonstrate the performance of our electronics with the Teststand for an Overpressurised Argon Detector (TOAD) at Fermilab in the US at a range of pressures and gas mixtures up to , reading out channels from a Multi-Wire Proportional Chamber (MWPC). The operation took place between April and July of 2024. We measure the noise characteristics of the system to be sufficiently low, and we identify sources of noise that can be further mitigated in the next iteration. We also note that the cooling scheme used in the test requires improvement before full-scale deployment. Despite these necessary improvements, we show that the system can fulfil the needs of a HPgTPC for a fraction of the price of collider readout electronics.
@@ 1 Introduction
The detector’s position resolution is a critical factor in achieving the required momentum resolution and detection thresholds for complex low-energy interactions with high multiplicity. The resolution obtainable is closely related to the number of readout channels that the detector can accommodate and the sampling frequency of those channels. Whilst the numbers are not final, instrumenting a diameter TPC sufficiently to achieve the desired low energy thresholds for the DUNE would require channels with sampling frequency [1]. The development of readout electronics in the particle physics community has previously focused on the design and operation of high-throughput radiation-hard electronics required at the Large Hadron Collider (LHC)-based experiments [4]. Readout electronics from sPHENIX [5] or the TPC of A Large Ion Collider Experiment (ALICE) [6] require a high-specification Field Programmable Gate Array (FPGA) to readout groups of 10 front-end cards. Naively scaling such designs for neutrino experiments would be prohibitively expensive. This is due to the hundreds of thousands of channels required to meet DUNE’s resolution needs. Hence, this approach would result in a total cost larger than possible, given budgetary constraints for the DUNE near detector. In fact, the electronics were previously listed as the most expensive component for the Gaseous Argon Near Detector (ND-GAr) for DUNE.
@@
In the following sections we present the design of a scalable and highly re-configurable FPGA-based solution optimised to be affordable for the ND-GAr DUNE Phase-II detector [3]. This system delivers significant cost-saving compared to the systems described above while showing no performance degradation in an HPgTPC environment. A prototype has been produced and was then sent for testing to Fermilab as part of the Teststand for an Overpressurised Argon Detector (TOAD). Whilst DUNE has not decided on an amplification stage for its HPgTPC, a Gas Electron Multiplier (GEM)-based solution is the current frontrunner. However, TOAD used a Multi-Wire Proportional Chamber (MWPC) from ALICE, due to its availability at the time of the test.
2 TOAD
The TOAD detector design and operation largely follow that of the UK high-pressure platform, described in [7]. The detector was sent to Fermilab, where it was commissioned, and the newly developed readout electronics were installed. These readout electronics interface with an ALICE Outer Read-Out Chamber (OROC), meaning TOAD was a full slice test of a HPgTPC.
2.1 ALICE Multi-Wire Proportional Chambers
Given that the ALICE TPC [8, 6] required an upgrade [9], the Read-Out Chambers can be repurposed. The OROC used in TOAD was part of the ALICE ‘test’ ROCs, rather than a production ROC. The specifications of the OROC used are detailed at length in [7].
The ROCs employ a typical wire plane schematic: an anode-wire grid above the pad plane, a cathode-wire grid and a gating grid to separate the drift volume from the readout. The pad plane and the three wire planes are all spaced apart. All wires run in the azimuthal direction (approximately vertically in the orientation we have the chamber) and have different separations in the various planes. The length of the shorter and longer parallel sides of the OROC are and , respectively. The perpendicular length of the OROC is . Voltages are applied to the anode and gating grid wires, while the cathode wires are held at ground. The pad plane consists of 9984 pads in total, of two sizes, which are summarised in table˜1. Here, “inner” refers to the rows closer to the shorter parallel side, whereas “outer” refers to rows closer to the longer side of the OROC. The difference in pad dimensions is due to the trapezoidal shape of the ROC. The readout electronics must interface with this ROC. Therefore, groups of 21 or 22 pads are read out via a socket on the back of the OROC. With the sockets grouped in threes, our Front-End Cards connect directly to these to read out a total of 64 channels each.
| Pad Dimensions (mm2) | Pad Rows | Number of Pads | |
| Outer Pads | 6 15 | 32 | 4032 |
| Inner Pads | 6 10 | 64 | 5952 |
| Total | - | 96 | 9984 |
2.2 Pressure vessel
TOAD utilises the same pressure vessel described in [7, 10], which allows the testing of the readout electronics with an ALICE OROC at pressures up to (). The vessel has an inner diameter of and an outer diameter of , made from Stainless Steel. The vessel has domed ends, giving a total volume of . The vessel can be opened by removing one of the domed faces. An image of the vessel in the test environment is shown in figure˜1. Further details of the vessel properties can be found in [7, 10]. The way in which the feedthroughs were utilised for TOAD differs from the original setup detailed in [7]. In addition, the vacuum system was replaced. There are various KF25 and KF40 flanges on the vessel, which are used for the High Voltage (HV), gas, vacuum and readout-electronics systems feedthroughs. The gas and evacuation system is connected to the vessel so that the vessel can be evacuated to to ensure the purity of gas in the detector. The pump system was upgraded to an Edwards Turbo Cart, consisting of a backing pump, a turbo pump and a controller. The gas system allowed the vessel to be filled with up to of gas mixtures, with the gas mix controlled by filling to a given partial pressure of gas from up to four different gas cylinders connected to the system. For the run described in this paper, three input gases were mixed: Ar-CH4 with a molar ratio of 92:8, pure Ar, and pure N2.
The TPC is the same as that described in [7]. It utilises a HV drift cathode, a field cage, a terminating plate, and the OROC. There are three HV feedthroughs on the vessel, which are used to apply voltages to the drift cathode, anode and gating grid. Other feedthroughs are also used to ensure the equipment inside the vessel is all safely grounded. The cathode and OROC anode were previously tested with voltages up to and respectively, with the maximum achievable voltage varying with the gas mixture used [7]. The results shown in this paper use various configurations of HV, gas mixtures and pressures. These are summarised in table˜2. The measurements outlined in this paper are noise characterisation studies, and so the particle tracking ability of the TPC is not employed.
| 1 | Air | 1.0 | 0 | 0 | 0 |
| 2 | Ar:CH4, 96:4 | 4.5 | 0 | 0 | 0 |
| 3 | Ar:CH4, 98:2 | 3.0 | 1500 | 0 | |
| 4 | Ar:CH4, 98:2 | 3.0 | 1500 |
In addition to the connections described above, the readout electronics were powered and streamed data through KF40 feedthroughs on the vessel door. One feedthrough was used to provide power, and three RJ45 feedthroughs are used to control, configure, monitor and stream data to the readout electronics. These were connected directly to a Commercial Off the Shelf (COTS) network switch, which was, in turn, connected to two DELL PowerEdge servers: one running CCM software, and the other Data Acquisition (DAQ) software – both of which are outlined in detail in section˜3.3.
@@ 3 Readout setup
A key challenge for future neutrino experiments is the harsh conditions in which the electronics must operate. In contrast to collider experiments, those conditions are not due to extensive radiation but instead factors such as cryogenic temperatures present in Liquid Argon (LAr)-based detectors and high pressure, such as in an HPgTPC based detector. In the case of the high-pressure detectors, an important consideration is the number of feedthroughs required to power and communicate with the electronics, so there is a need to aggregate data inside the vessel. Whilst power dissipation and thermal management present additional challenges, they can be addressed using water cooling through a small number of vessel penetrations. Hence, a readout system will need to optimise the number of cables required to stream data from the large number of analogue readout channels. In addition, based on the physics requirements for the DUNE near detector (ND) complex outlined in the Conceptual Design Report [1], the proposed readout system should not constrain or limit the advantages provided by the installation of a magnetised high-pressure gaseous TPC in the complex. Such requirements foresee enhanced tracking capabilities in comparison to the LAr-based detector counterpart for more accurate event classification. A summary of the requirements taken into account can be seen in table˜3 with the equivalent design choice and the technology implementation. Obviously, choices like the number of links aggregated into a single FPGA are driven based on the capabilities of the package as well as the cost, making it an optimisation process based on physics versus cost.
| Requirement | Design choice | Implementation |
| Detection threshold ((MeV) protons) | High-Pressure TPC | Reduced feedthroughs |
| Track separation () | Analogue to Digital Converter (ADC) sampling | |
| Num. of detector channels | (700k) | |
| Low event rates () | Gigabit readout | Kintex Ultrascale FPGA |
@@
The proposed solution outlined in this paper uses a fast-sampling FEC with zero-suppression capability and a custom FPGA-based Power, Aggregation and Timing (PAT) card, both residing in-situ within the high-pressure volume. Data are then streamed over a standard Ethernet-based network to the DUNE DAQ software operating on standard servers, avoiding extensive requirements within the pressurised environment. The block diagram of the proposed readout system is shown in figure˜2.
@@
Compared to a collider experiment, the channels in neutrino experiments have very low occupancy (in the similarly sized ALICE, collisions occur at and each generate thousands of tracks, whereas in DUNE, beam spills happen at and generate only tens of tracks per spill). This difference means far less processing is required as there are far fewer above-threshold hits. Therefore, the system can have fewer FPGAs per channel (1 per channels with the particular FECs in this design) and use much lower specification FPGAs as we stream out all hits to a software-based DAQ. Comparisons with the ALICE and sPHENIX readout electronics are shown in table˜4.
| System | Event Rate | Sampling Rate | Channels/board |
| sPHENIX | 192 | ||
| ALICE | Vary per detector region | ||
| ND-GAr | 3584 |
Furthermore, the lower data rates from the FEC allow the native transceivers on the ADC Application Specific Integrated Circuits to be used to communicate with the PAT, so we also do not need the additional transceiver modules that are required for higher occupancy detectors [9]. These differences result in a much lower cost to instrument the hundreds of thousands of channels necessary for a neutrino HPgTPC.
@@ 3.1 Hardware layout
As described, our system requires two types of custom hardware boards. The first type are the FEC boards, which host the digitiser ASICs and connect to the detector itself. The second type are the PAT boards, which utilise an FPGA device for interfacing and aggregating the data for a multitude of FECs as well as providing them with power and distributing clock signals. Both the FECs and the PATs are custom-designed, with the FECs designed for the TOAD test beam detector and its ALICE OROC. The flexibility of the FPGA-based design of the PAT cards means that the PAT can be reused even if the FECs change, so, whilst designed for TOAD, this board could be used with minimal modifications for the final ND-GAr detector or other detectors with similar data volume and occupancy. In designing the hardware, we avoided components with known air gaps, e.g. lidded FPGA packages. However, as there is no standard for electronics operation in the regime, a goal of TOAD is to test the operation of components at high pressures.
@@ 3.1.1 Front-End Card
GEMs are currently the leading candidate for the charge readout for ND-GAr. The SAMPA ASIC used for this readout system was developed by the ALICE collaboration for the GEM upgrade of their TPC [9]. This upgraded detector has similar characteristics to those needed for ND-GAr, so SAMPA is expected to meet the requirements of ND-GAr. The parameters of the SAMPA measured by ALICE can be seen in table˜5 [11, 9]. The electronics system described here is, therefore, designed around deploying the SAMPA ASIC [11] used by ALICE, in a high-pressure environment, and at a much lower cost than in the full ALICE system.
@@
Another important reason for the choice of the SAMPA chips is their ability to sample at with fast ADCs. Gas detectors have about an order of magnitude faster electron drift velocities than liquid; hence, this high sample frequency is necessary.
For TOAD, the ALICE MWPC has a group of three connectors per slot, which correspond to a total of 64 channels. Given that the SAMPA ASIC has only 32 channels, the FEC designed for TOAD, which can be seen in figure˜3, has three connectors and hosts two SAMPAs to capture all 64 channels. The final ND-GAr amplification stages may have a different number of channels per connector; however, the more SAMPAs per FEC, the lower the cost of the system, so designing for two SAMPAs represents a reasonable worst-case cost. To minimise the required readout channels connecting to the PAT board, the two SAMPAs on each FEC are daisy-chained, as the low occupancy of neutrino experiments does not require the full bandwidth of the SAMPA readout link. However, as this is a prototype detector, the layout of the FEC has been designed to allow each SAMPA to be read out individually with minimal changes to the fitted components and no re-design of the Printed Circuit Board (PCB).
For TOAD, the FECs were operated in the configuration shown in table˜6. This configuration gives an expected current draw of /channel with the FEC operating at full capacity. It is, however, not expected that the FEC will have to operate at full power due to low occupancy and zero suppression. Whilst the design of the FEC is specific to TOAD, its operation at TOAD showcases the expected performance of SAMPA more generally in the type of environment that is required for HPgTPCs, which has not been previously tested.
@@ 3.1.2 Power Aggregation and Timing cards
The PAT card’s primary function is to interface between the FEC cards and a standard Ethernet network, where data can be streamed to servers for data selection and storage. The prototype, designed at Imperial College London, utilises the advantages of FPGAs as the main data processing unit for the PAT. An image of the prototype board can be seen in figure˜4.
Data aggregation is a primary requirement that drives the design of the PAT. The cards are designed to take inputs from FECs each streamed over a differential pair into the FPGA package. The received data payloads from the many FECs are then extracted and propagated through a total of buffers using Round-Robin (RR) logic to aggregate them into a single output stream. This output data stream is then wrapped into an Ethernet-based protocol and sent over the FPGA serial transceivers to the COTS network switch. The full layout can be seen in figure˜5. Due to the low occupancy of the detector, the aggregation factor of 56 to 1 can be achieved easily within the FPGA fabric.
As well as data aggregation, the powering and timing roles of the PAT are also crucial. For powering, we investigated two possible solutions, for which a summary of key metrics is provided in table˜7. The first solution provides to each PAT via dedicated terminals on the board. The FPGA DC/DC converter then outputs which is further fed into the FEC DC/DC converter to be transformed to the needed by the FEC on-board Low-Dropout regulator (LDO). The advantage of this scheme is that each PAT card receives a dedicated power line, minimising potential interference with signal traces on the board. The disadvantage is that ND-GAr will need dedicated power feedthroughs in the pressure vessel, which increases the overall feedthrough count.
The second ‘power over serial’ powering scheme is inspired by Power over Ethernet (PoE), and aims to allow PAT cards placed within the pressure vessel to receive power over the incoming network cable. The circuit diagram used can be seen in figure˜6. However, some modifications to the normal PoE standard are required due to the intended environment. PoE works by injecting a Direct Current (DC) component on top of the alternating data component using transformers. Copying the same methodology for a magnetised detector wouldn’t be viable, as transformer performance may be impacted by the surrounding magnetic field. Therefore, an alternative system was developed using air-cored inductors to extract the differential signal from the bias voltage.
@@
The implementation on an early prototype for the PAT cards was found not to give sufficient signal stability, so further optimisations are required to achieve a production-level system with this powering option. This work is underway with a new dedicated hardware board allowing modifications to both the power injection and the power decoupling circuits for further analysis. However, as it has not yet converged, we therefore chose the first solution (direct power) for the board design discussed here.
In addition, a study, summarised in table˜7, was done to estimate the number of feedthroughs needed for both options and also a third potential solution which would use optical fibres to stream out the data and a separate power line for each PAT. This option allows several PAT cards to share one output feedthrough as optical fibres are smaller in form factor than the electrical connectors needed.
@@
Decisions on the power distribution of the system also have an impact on the timing. For the Power over Serial method, timing can be provided to the upstream PAT card that is outside the pressure vessel and all the cards within the vessel can then be synchronised by the serial transceiver recovered clock. In the case of Direct Power, the data stream is separate from power and is running through commercial Small Form-Factor Pluggable (SFP) modules. We therefore require a separate dedicated input for clock distribution. The PAT cards are designed to be able to interface with the DUNE timing system [12] over an optical fibre. In order to implement this, extra optical feedthroughs are required to receive all the timing fibres into the pressure vessel. The number of feedthroughs can be mitigated by fanning out the signal inside the vessel, but there are limits to the number of boards that can be supplied from any one fibre. In both cases, the clock received at the PAT cards in the pressure vessel is transmitted from the FPGA to a dedicated clock synthesiser for cleaning and then further transmitted to the FECs.
Furthermore, distribution of the FEC clock, synchronisation, and slow control signals happen, depending on the use case, via fan-out chips or multiplexers. Given that the designed readout is synchronous to the DUNE clock, all the FEC cards are expected to receive the same copy of the clock synchronisation signal. For this reason, an organisation of eight FECs in seven groups has been designed. The advantage of this method relies on the layout design of the PCB. The chosen FPGA can drive seven individual groups with the same clock synchronisation signal pulse. Each of these seven individual groups is then copied in phase to all eight FECs connected to that pin. The main reason for deciding to fan-out 1:8 was to ensure that the clock quality provided to all the FECs wasn’t impacted by the fan-out ASIC. Equivalently, for the slow control interface, a series of multiplexers is implemented, allowing individual access to all the FECs, providing enhanced telemetry of the FEC operation.
Finally, the PAT cards are mounted at three positions on the OROC, in such a way that when 18 OROCs are mounted into a circle, the PAT cards are shared between neighbouring OROCs. In this way, the cable lengths required between FECs and PAT cards are minimised, and so, these connections are optimised. The PAT cards are mounted on aluminium plates, which, through thermal contact with the ROC, provide some amount of passive cooling. An image of the FECs and PATs connected to the OROC inside the vessel is shown in figure˜7.